\begin{thebibliography}{25}
\providecommand{\natexlab}[1]{#1}

\bibitem[{Bajaj et~al.(2018)Bajaj, Campos, Craswell, Deng, Gao, Liu, Majumder,
  McNamara, Mitra, Nguyen, Rosenberg, Song, Stoica, Tiwary, and
  Wang}]{bajaj2018msmarcohumangenerated}
Payal Bajaj, Daniel Campos, Nick Craswell, Li~Deng, Jianfeng Gao, Xiaodong Liu,
  Rangan Majumder, Andrew McNamara, Bhaskar Mitra, Tri Nguyen, Mir Rosenberg,
  Xia Song, Alina Stoica, Saurabh Tiwary, and Tong Wang. 2018.
\newblock \href {https://arxiv.org/abs/1611.09268} {Ms marco: A human generated
  machine reading comprehension dataset}.
\newblock \emph{Preprint}, arXiv:1611.09268.

\bibitem[{Craswell et~al.(2021)Craswell, Mitra, Yilmaz, Campos, and
  Lin}]{craswell2021msmarcobenchmarkingranking}
Nick Craswell, Bhaskar Mitra, Emine Yilmaz, Daniel Campos, and Jimmy Lin. 2021.
\newblock \href {https://arxiv.org/abs/2105.04021} {Ms marco: Benchmarking
  ranking models in the large-data regime}.
\newblock \emph{Preprint}, arXiv:2105.04021.

\bibitem[{Devlin et~al.(2019)Devlin, Chang, Lee, and
  Toutanova}]{devlin-etal-2019-bert}
Jacob Devlin, Ming-Wei Chang, Kenton Lee, and Kristina Toutanova. 2019.
\newblock \href {https://doi.org/10.18653/v1/N19-1423} {{BERT}: Pre-training of
  deep bidirectional transformers for language understanding}.
\newblock In \emph{Proceedings of the 2019 Conference of the North {A}merican
  Chapter of the Association for Computational Linguistics: Human Language
  Technologies, Volume 1 (Long and Short Papers)}, pages 4171--4186,
  Minneapolis, Minnesota. Association for Computational Linguistics.

\bibitem[{Douze et~al.(2025)Douze, Guzhva, Deng, Johnson, Szilvasy, Mazaré,
  Lomeli, Hosseini, and Jégou}]{douze2025faisslibrary}
Matthijs Douze, Alexandr Guzhva, Chengqi Deng, Jeff Johnson, Gergely Szilvasy,
  Pierre-Emmanuel Mazaré, Maria Lomeli, Lucas Hosseini, and Hervé Jégou.
  2025.
\newblock \href {https://arxiv.org/abs/2401.08281} {The faiss library}.
\newblock \emph{Preprint}, arXiv:2401.08281.

\bibitem[{Formal et~al.(2021{\natexlab{a}})Formal, Lassance, Piwowarski, and
  Clinchant}]{formal2021spladev2sparselexical}
Thibault Formal, Carlos Lassance, Benjamin Piwowarski, and Stéphane Clinchant.
  2021{\natexlab{a}}.
\newblock \href {https://arxiv.org/abs/2109.10086} {Splade v2: Sparse lexical
  and expansion model for information retrieval}.
\newblock \emph{Preprint}, arXiv:2109.10086.

\bibitem[{Formal et~al.(2021{\natexlab{b}})Formal, Piwowarski, and
  Clinchant}]{formal2021spladesparselexicalexpansion}
Thibault Formal, Benjamin Piwowarski, and Stéphane Clinchant.
  2021{\natexlab{b}}.
\newblock \href {https://arxiv.org/abs/2107.05720} {Splade: Sparse lexical and
  expansion model for first stage ranking}.
\newblock \emph{Preprint}, arXiv:2107.05720.

\bibitem[{Hofstätter et~al.(2021)Hofstätter, Althammer, Schröder, Sertkan,
  and Hanbury}]{hofstätter2021improvingefficientneuralranking}
Sebastian Hofstätter, Sophia Althammer, Michael Schröder, Mete Sertkan, and
  Allan Hanbury. 2021.
\newblock \href {https://arxiv.org/abs/2010.02666} {Improving efficient neural
  ranking models with cross-architecture knowledge distillation}.
\newblock \emph{Preprint}, arXiv:2010.02666.

\bibitem[{Jiang et~al.(2023)Jiang, Xu, Gao, Sun, Liu, Dwivedi-Yu, Yang, Callan,
  and Neubig}]{jiang-etal-2023-active}
Zhengbao Jiang, Frank Xu, Luyu Gao, Zhiqing Sun, Qian Liu, Jane Dwivedi-Yu,
  Yiming Yang, Jamie Callan, and Graham Neubig. 2023.
\newblock \href {https://doi.org/10.18653/v1/2023.emnlp-main.495} {Active
  retrieval augmented generation}.
\newblock In \emph{Proceedings of the 2023 Conference on Empirical Methods in
  Natural Language Processing}, pages 7969--7992, Singapore. Association for
  Computational Linguistics.

\bibitem[{Karpukhin et~al.(2020)Karpukhin, Oguz, Min, Lewis, Wu, Edunov, Chen,
  and Yih}]{karpukhin-etal-2020-dense}
Vladimir Karpukhin, Barlas Oguz, Sewon Min, Patrick Lewis, Ledell Wu, Sergey
  Edunov, Danqi Chen, and Wen-tau Yih. 2020.
\newblock \href {https://doi.org/10.18653/v1/2020.emnlp-main.550} {Dense
  passage retrieval for open-domain question answering}.
\newblock In \emph{Proceedings of the 2020 Conference on Empirical Methods in
  Natural Language Processing (EMNLP)}, pages 6769--6781, Online. Association
  for Computational Linguistics.

\bibitem[{Khattab and
  Zaharia(2020)}]{khattab2020colbertefficienteffectivepassage}
Omar Khattab and Matei Zaharia. 2020.
\newblock \href {https://arxiv.org/abs/2004.12832} {Colbert: Efficient and
  effective passage search via contextualized late interaction over bert}.
\newblock \emph{Preprint}, arXiv:2004.12832.

\bibitem[{Lee et~al.(2019)Lee, Chang, and Toutanova}]{lee-etal-2019-latent}
Kenton Lee, Ming-Wei Chang, and Kristina Toutanova. 2019.
\newblock \href {https://doi.org/10.18653/v1/P19-1612} {Latent retrieval for
  weakly supervised open domain question answering}.
\newblock In \emph{Proceedings of the 57th Annual Meeting of the Association
  for Computational Linguistics}, pages 6086--6096, Florence, Italy.
  Association for Computational Linguistics.

\bibitem[{Lewis et~al.(2021)Lewis, Perez, Piktus, Petroni, Karpukhin, Goyal,
  Küttler, Lewis, tau Yih, Rocktäschel, Riedel, and
  Kiela}]{lewis2021retrievalaugmentedgenerationknowledgeintensivenlp}
Patrick Lewis, Ethan Perez, Aleksandra Piktus, Fabio Petroni, Vladimir
  Karpukhin, Naman Goyal, Heinrich Küttler, Mike Lewis, Wen tau Yih, Tim
  Rocktäschel, Sebastian Riedel, and Douwe Kiela. 2021.
\newblock \href {https://arxiv.org/abs/2005.11401} {Retrieval-augmented
  generation for knowledge-intensive nlp tasks}.
\newblock \emph{Preprint}, arXiv:2005.11401.

\bibitem[{Mikolov et~al.(2013)Mikolov, Chen, Corrado, and
  Dean}]{mikolov2013efficientestimationwordrepresentations}
Tomas Mikolov, Kai Chen, Greg Corrado, and Jeffrey Dean. 2013.
\newblock \href {https://arxiv.org/abs/1301.3781} {Efficient estimation of word
  representations in vector space}.
\newblock \emph{Preprint}, arXiv:1301.3781.

\bibitem[{M{\"u}ller and Gurevych(2009)}]{muller-gurevych-2009-study}
Christof M{\"u}ller and Iryna Gurevych. 2009.
\newblock \href {https://aclanthology.org/D09-1139/} {A study on the semantic
  relatedness of query and document terms in information retrieval}.
\newblock In \emph{Proceedings of the 2009 Conference on Empirical Methods in
  Natural Language Processing}, pages 1338--1347, Singapore. Association for
  Computational Linguistics.

\bibitem[{Ni et~al.(2022)Ni, Qu, Lu, Dai, Hernandez~Abrego, Ma, Zhao, Luan,
  Hall, Chang, and Yang}]{ni-etal-2022-large}
Jianmo Ni, Chen Qu, Jing Lu, Zhuyun Dai, Gustavo Hernandez~Abrego, Ji~Ma,
  Vincent Zhao, Yi~Luan, Keith Hall, Ming-Wei Chang, and Yinfei Yang. 2022.
\newblock \href {https://doi.org/10.18653/v1/2022.emnlp-main.669} {Large dual
  encoders are generalizable retrievers}.
\newblock In \emph{Proceedings of the 2022 Conference on Empirical Methods in
  Natural Language Processing}, pages 9844--9855, Abu Dhabi, United Arab
  Emirates. Association for Computational Linguistics.

\bibitem[{Nogueira et~al.(2020)Nogueira, Jiang, Pradeep, and
  Lin}]{nogueira-etal-2020-document}
Rodrigo Nogueira, Zhiying Jiang, Ronak Pradeep, and Jimmy Lin. 2020.
\newblock \href {https://doi.org/10.18653/v1/2020.findings-emnlp.63} {Document
  ranking with a pretrained sequence-to-sequence model}.
\newblock In \emph{Findings of the Association for Computational Linguistics:
  EMNLP 2020}, pages 708--718, Online. Association for Computational
  Linguistics.

\bibitem[{Raffel et~al.(2023)Raffel, Shazeer, Roberts, Lee, Narang, Matena,
  Zhou, Li, and Liu}]{raffel2023exploringlimitstransferlearning}
Colin Raffel, Noam Shazeer, Adam Roberts, Katherine Lee, Sharan Narang, Michael
  Matena, Yanqi Zhou, Wei Li, and Peter~J. Liu. 2023.
\newblock \href {https://arxiv.org/abs/1910.10683} {Exploring the limits of
  transfer learning with a unified text-to-text transformer}.
\newblock \emph{Preprint}, arXiv:1910.10683.

\bibitem[{Robertson et~al.(1994)Robertson, Walker, Jones, Hancock-Beaulieu, and
  Gatford}]{conf/trec/RobertsonWJHG94}
Stephen~E. Robertson, Steve Walker, Susan Jones, Micheline Hancock-Beaulieu,
  and Mike Gatford. 1994.
\newblock \href
  {http://dblp.uni-trier.de/db/conf/trec/trec94.html#RobertsonWJHG94} {Okapi at
  trec-3.}
\newblock In \emph{TREC}, volume 500-225 of \emph{NIST Special Publication},
  pages 109--126. National Institute of Standards and Technology (NIST).

\bibitem[{Salton and Buckley(1988)}]{salton_termweighting_1988}
Gerard Salton and Christopher Buckley. 1988.
\newblock \href {https://doi.org/10.1016/0306-4573(88)90021-0} {Term-weighting
  approaches in automatic text retrieval}.
\newblock \emph{Information Processing \& Management}, 24(5):513--523.

\bibitem[{Sanh et~al.(2020)Sanh, Debut, Chaumond, and
  Wolf}]{sanh2020distilbertdistilledversionbert}
Victor Sanh, Lysandre Debut, Julien Chaumond, and Thomas Wolf. 2020.
\newblock \href {https://arxiv.org/abs/1910.01108} {Distilbert, a distilled
  version of bert: smaller, faster, cheaper and lighter}.
\newblock \emph{Preprint}, arXiv:1910.01108.

\bibitem[{Santhanam et~al.(2022)Santhanam, Khattab, Saad-Falcon, Potts, and
  Zaharia}]{santhanam-etal-2022-colbertv2}
Keshav Santhanam, Omar Khattab, Jon Saad-Falcon, Christopher Potts, and Matei
  Zaharia. 2022.
\newblock \href {https://doi.org/10.18653/v1/2022.naacl-main.272}
  {{C}ol{BERT}v2: Effective and efficient retrieval via lightweight late
  interaction}.
\newblock In \emph{Proceedings of the 2022 Conference of the North American
  Chapter of the Association for Computational Linguistics: Human Language
  Technologies}, pages 3715--3734, Seattle, United States. Association for
  Computational Linguistics.

\bibitem[{Thakur et~al.(2021)Thakur, Reimers, Rücklé, Srivastava, and
  Gurevych}]{thakur2021beirheterogenousbenchmarkzeroshot}
Nandan Thakur, Nils Reimers, Andreas Rücklé, Abhishek Srivastava, and Iryna
  Gurevych. 2021.
\newblock \href {https://arxiv.org/abs/2104.08663} {Beir: A heterogenous
  benchmark for zero-shot evaluation of information retrieval models}.
\newblock \emph{Preprint}, arXiv:2104.08663.

\bibitem[{Wang et~al.(2024)Wang, Huang, Tu, Wang, Huang, Laskar, and
  Bhuiyan}]{wang2024utilizingbertinformationretrieval}
Jiajia Wang, Jimmy~X. Huang, Xinhui Tu, Junmei Wang, Angela~J. Huang,
  Md~Tahmid~Rahman Laskar, and Amran Bhuiyan. 2024.
\newblock \href {https://arxiv.org/abs/2403.00784} {Utilizing bert for
  information retrieval: Survey, applications, resources, and challenges}.
\newblock \emph{Preprint}, arXiv:2403.00784.

\bibitem[{Yang et~al.(2024)Yang, Jiang, and
  Baldwin}]{yang-etal-2024-language-bias}
Jinrui Yang, Fan Jiang, and Timothy Baldwin. 2024.
\newblock \href {https://doi.org/10.18653/v1/2024.mrl-1.23} {Language bias in
  multilingual information retrieval: The nature of the beast and mitigation
  methods}.
\newblock In \emph{Proceedings of the Fourth Workshop on Multilingual
  Representation Learning (MRL 2024)}, pages 280--292, Miami, Florida, USA.
  Association for Computational Linguistics.

\bibitem[{Yates et~al.(2021)Yates, Nogueira, and
  Lin}]{yates-etal-2021-pretrained}
Andrew Yates, Rodrigo Nogueira, and Jimmy Lin. 2021.
\newblock \href {https://doi.org/10.18653/v1/2021.naacl-tutorials.1}
  {Pretrained transformers for text ranking: {BERT} and beyond}.
\newblock In \emph{Proceedings of the 2021 Conference of the North American
  Chapter of the Association for Computational Linguistics: Human Language
  Technologies: Tutorials}, pages 1--4, Online. Association for Computational
  Linguistics.

\end{thebibliography}

\pdfoutput=1

\documentclass[11pt]{article}

\usepackage[final]{acl}

\usepackage{times}
\usepackage{latexsym}
\usepackage{amsmath}

\usepackage[T1]{fontenc}

\usepackage[utf8]{inputenc}

\usepackage{microtype}

\usepackage{inconsolata}

\usepackage{graphicx}

\title{Semantic Search for Information Retrieval}

\author{
  Kayla Farivar \\
  University of California, Berkeley \\
  \texttt{kaylafari@berkeley.edu}
}

\begin{document}
\maketitle
\begin{abstract}
Information retrieval systems have progressed notably from lexical techniques such as BM25 and TF-IDF to modern semantic retrievers. This survey provides a brief overview of the BM25 baseline, then discusses the architecture of modern state-of-the-art semantic retrievers. Advancing from BERT, we introduce dense bi-encoders (DPR), late-interaction models (ColBERT), and neural sparse retrieval (SPLADE). Finally, we examine MonoT5, a cross-encoder model. We conclude with common evaluation tactics, pressing challenges, and propositions for future directions.
\end{abstract}

\section{Introduction to Information Retrieval}
Information retrieval (IR) refers to returning relevant information based on a human-made query and a knowledge base to search (generally a large collection of documents). After taking a source of knowledge and a pre-made query, an IR system uses a retriever to return relevant information from the knowledge base. A retriever picks which information to return based on the semantic similarity of the query and chunks of each document \cite{muller-gurevych-2009-study}. Semantic similarity measures the closeness of the contextual meaning of two pieces of text. This is important due to the volatility of language - a small change in a sentence can completely shift the meaning of the sentence. IR is used for various applications. A search engine may use a reranking layer to give a user multiple options of what documents to look at, while another system may use the information to generate an answer from an LLM \cite{jiang-etal-2023-active}.

\section{History}
BM25 (Best Match 25) \cite{conf/trec/RobertsonWJHG94} is a purely lexical scoring function popularly used for ranking in information retrieval systems. Although this function used no semantic properties, it’s still highly regarded and used as a baseline against modern-day retrievers. BM25 is highly inspired by another formula for ranking called TF-IDF (Term Frequency - Inverse Document Frequency). Both translate documents into a Bag-of-words vector. Bag-of-words represents documents as a vector whose dimension size is of an all-encompassing set of vocabulary, each entry holding the word's frequency in a specific document (or zero if the term is absent) \citep{salton_termweighting_1988}. The returning vector explains the meaning of a document through the terms that could have been mentioned and the terms that had been mentioned. BM-25 deviates by introducing diminishing returns for term frequency, so as a term appears more and more often, it doesn’t increase how much it matters. Additionally, it standardizes documents so they aren’t awarded for being longer. 

In the early 2010s, semantic embeddings were popularized with the introduction of word2vec \citep{mikolov2013efficientestimationwordrepresentations}. Word2vec uses local context to determine similarity. Assume if we often saw sentences about lions, we could replace the word lion with the word giraffe. Word2vec would then correctly assume lions and giraffes have a close relationship as animals. BM25 kept steady as the strongest scoring function through the innovations around embeddings, the most famous being BERT \citep{devlin-etal-2019-bert}, until ORQA \cite{lee-etal-2019-latent} made a dense retriever that was very competitive with BM25’s scores. ORQA’s designs utilized BERT, a bidirectional encoder that provides highly contextual embedding outputs, and provided high motivation for semantic search \cite{yates-etal-2021-pretrained}.

\section{Dense Bi-Encoder Models}
\citet{karpukhin-etal-2020-dense} introduced one of the first Dense Bi-Encoders inspired by ORQA’s designs. The model uses BERT as an encoder and finds the top-k most similar queries to document pairs with a near-real-time algorithm from FAISS \citep{douze2025faisslibrary}, which uses complex index structures to speed up algorithms. With BERT, the last token in a set of embeddings is the [CLS] token, which includes an understanding of every token before it. They chose their formula for similarity as the dot product between the [CLS] tokens for the query and document. The bulk of their work has to do with their chosen method of negative training. Within training, there is one correct passage for a given query, negative training refers to the samples given to the model that are not correct. They tested a few collections of negative samples. First, they tried a series of three collections: random incorrect passages from their data, then only passages that BM25 retrieved but weren't correct, and lastly passages that were only correct for other sample questions. These sets of negative samples comparatively had little to no impact on their top-k accuracy. Next, they tried only using passages that were correct for other sample questions in the same batch during training and found a large improvement in their results. Further trialing found that adding one of BM25’s hard negatives (passages BM25 rated highly but were incorrect) to their samples improved their accuracy further (and adding more than one sample made little to no difference). The resulting model generalized well (small decrease in accuracy when trying independent sources), and its improved design significantly beat BM25 with a top-5 accuracy of 65.2

\section{Late-Interaction Models}
ColBERT \citep{khattab2020colbertefficienteffectivepassage} is a late-interaction model that gets an embedding per token for the input text. ColBERT adds an interesting feature in training called query augmentation. This adds padding to queries with the special [MASK] token. In BERT, this token is usually used for masking words in training texts, and by predicting the masked words, it helps BERT learn what words mean through context. With ColBERT, this additional padding allows the model to imagine what words might exist in the query around it. For example, if a query included the terms “American Independence Day,” it might imagine similar terms like “Fourth of July” or “U.S national holiday” to add near it. This allows for the query to be stronger and find more passages similar to it. For the separate document encoder, there is no padding, but as usual, the text is tokenized and an embedding is made for each resulting token. With two sets of query and document embeddings, a score is computed by finding the maximum value dot product of each query token’s embedding with any embedding from the document’s tokens. In ColBERT \cite{santhanam-etal-2022-colbertv2}, they speed up this process and add token weighting to the query to show stronger importance to select tokens. ColBERT's scoring allows for a high precision because one vector doesn't collect as many nuances and details, although it can be memory-intensive due to the count of tokens in queries and documents. It’s recommended in practice if the amount of documents that need processing is high, ColBERT should be used as a top-k reranking system. The idea is to use a primary retriever like BM25 to get the top-k documents and then feed that smaller k documents into ColBERT.

\section{Neural Sparse Retrieval}
SPLADE \cite{formal2021spladesparselexicalexpansion} is a neural sparse retriever that wanted to take more inspiration from BM25 and move away from dense embeddings in DPR retrievers. BM25 used a very sparse bag-of-words vector. SPLADE wanted to bridge the interpretability of sparse embeddings while training with neural networks like DPR models. To achieve this, SPLADE gets embeddings with BERT (or more commonly at this age, DistilBERT \citep{sanh2020distilbertdistilledversionbert} due to its balance of efficiency and performance) for each token in the query, and it then reuses BERT's masked language modeling head to get logits over the entire vocabulary per token. Logits represent how likely each term is to appear there, and they give an idea of the importance of each term. SPLADE then uses a log saturation function to add up each of the logits into one score. The purpose of using log saturation $ \left( \sum (\log (1+w) \text{ where w = logit weight} \right) $  instead of a simple sum is to suppress large values. This process results in a vocabulary-length vector of various logits indicating its highest predicted relevance. ReLU is then used to remove all negatives and create the final sparse vector. The FLOPS regularizer was also introduced to penalize the loss for frequently active terms across documents. The same actions are repeated for the document with the same encoder, and the dot product of the two is used as their similarity score. 

A year after SPLADE's introduction, a few extra optimizations were introduced for SPLADE v2 \citep{formal2021spladev2sparselexical}. SPLADE v2 further enforces sparsity by updating its FLOPS regularizer to penalize the loss for having too many non-zero outputs. The rest of their optimizations were made into optional settings with a differently named model for each. Inspired by other models (including ColBERT), SPLADE-max uses max pooling to aggregate the logits rather than the previous log saturation sum. SPLADE-doc is a new encoder specifically for documents. When the query is encoded, the system may mark tokens that weren’t exactly in the query as relevant tokens and give each relevant token a weight. This allows for a more versatile search query. The new SPLADE-doc is similar to before, but only tokens that were in the document can have a weight, and all of the weights are set to be equal. This optimization was purely made for efficiency and had little effect on the accuracy. Lastly, they took a new approach called model distillation \citep{hofstätter2021improvingefficientneuralranking}. The idea is to take a highly accurate model (in this case, a cross-encoder) and train triplets (a query, a correct document, and an incorrect document), get this model to score new triplets, and train a lighter weight model (SPLADE in this case) to output these same scores. This allows for a teacher (a cross encoder) to “teach” the student (SPLADE). This new model is called DistilSPLADE-max.

\section{Cross-Encoder Models}
Information Retrieval was dominated by BERT pretrained models until 2020 \citep{wang2024utilizingbertinformationretrieval}, when \citet{nogueira-etal-2020-document} made the cross-encoder MonoT5. T5 \citep{raffel2023exploringlimitstransferlearning} is an encoder-decoder transformer, the encoder takes in a text input and the decoder is run autoregressively to produce a text output. Simplified, T5 pretrains very similarly to BERT's masking training, but with multiple tokens being masked. T5 uses special tokens like <x> <y> <z>, for example, and masks tokens; the model's job is then to predict what goes under each token. This helps the model learn the meaning of words in the context of a sentence, but also learn how to generate context. MonoT5 took in the sequence “Query: q Document: d Relevant: “ (where q and d were the respective documents) and outputed the sequence “true” or “false”. MonoT5 reframed information retrieval and helped popularize using “target words” (like true or false in this instance) to transform sequence-to-sequence models into acting similarly to most other models. To simply get the probabilities of each document, they apply softmax to the logits of the tokens “true” and “false”. This works very well because the target words “true” and “false” are only one token, which allows them to extract the probability of the document. The steps to finetune T5 into MonoT5, can be repeated with larger versions of t5 \citep{ni-etal-2022-large}, GPT 3.5, GPT 4.0, or other LLMS.

\section{Benchmarking and Evaluation}
The most popular evaluation dataset is MS MARCO (MiscroSoft MAchine Reading COmprehension) \citep{bajaj2018msmarcohumangenerated}, debuting in almost every modern paper on information retrieval. The expansive dataset scraped Bing to produce over a million anonymized questions, each with a human-generated answer, and almost 9 million passages. Their team also created a leaderboard with bootstrapping and private queries to validate performance. Although the MS MARCO leader board helps establish a history of improvement, other benchmarks like BEIR (Benchmarking-IR) \citep{thakur2021beirheterogenousbenchmarkzeroshot} should be used in practice to avoid leaderboard chasing and prove well-roundedness \citep{craswell2021msmarcobenchmarkingranking}.

\section{Conclusion and Future Directions}
In summary, semantic search retrievers overtook the lexical innovations of BM25 at their place in information retrieval. These innovations have allowed for the creation of real-time RAG-based \citep{lewis2021retrievalaugmentedgenerationknowledgeintensivenlp} virtual assistants, enhanced search engines, and stronger data analysis for personalized recommendations and real-time information access. While great progress has been made, there are significant challenges that persist. English-centric datasets overlook multilingual compatibility and low-resource languages \citep{yang-etal-2024-language-bias}. Further research is necessary to allow for the collective benefits of technological advancement.

\bibliography{custom, anthology}

\end{document}